OpenStudy (snowcrystal):

HElpppppp!!!! I don't get this stuff!! I will Fan and Medal!!! Use summation notation to write the series 2+4+6+8+... for 10 terms. In each term of these images, the lower limit of the summation notation is either "n=1" or "n=0". Answer choices will be in comments..

OpenStudy (snowcrystal):

\[\sum_{n=1}^{10}2n\] \[\sum_{n=1}^{10}(n+2)\] \[\sum_{n=0}^{10}n\] \[\sum_{n=0}^{10}2n\]

OpenStudy (snowcrystal):

@K12awesomeness

OpenStudy (k12awesomeness):

I don't know this, sorry..

OpenStudy (k12awesomeness):

I know someone who might though

OpenStudy (snowcrystal):

WHo

OpenStudy (k12awesomeness):

@Mudkillz ... he might know somethin

OpenStudy (k12awesomeness):

or @usercode3rror he may know somethin too

OpenStudy (snowcrystal):

i need all the help I can get with this stuff

OpenStudy (anonymous):

second term: 4 third term: 6 you're adding 2 each time and you're starting with 2 ------------------------------ so if you start with 2, and you add 2 say n times, then you're left with 2+2n this only works if you start with n = 0, if you want to start with n = 1, you would have to subtract 2 to get 2+2n-2 = 2n+0 = 2n So the general term is 2n

OpenStudy (anonymous):

there u go

OpenStudy (k12awesomeness):

So, that would be B, right or no?

OpenStudy (anonymous):

no its a

OpenStudy (k12awesomeness):

kk

OpenStudy (snowcrystal):

:3 thank you

OpenStudy (anonymous):

your welcome